Calculate the percentage composition by weight for a mixture of ammonium chloride, silicon dioxide and sodium chloride using the following data:

a.) After heating a mixture with a mass of 10 grams in a fume hood for 15 minutes. The mass of the sample decreased to 6.68 grams.

b.) You then add several 5 mL portions of water to the sample, stir for 2 minutes and collected the liquid in a beaker. After heating the silicon dioxide, cool in a desiccator, the mass was determined to be 3 grams.

% ammonium chloride _______ % silicon dioxide________ % sodium chloride _______
Answer #1

On heating the reaction occurs:

2 NaCl = 2 Na + Cl2

Where Cl2 is released in a gaseous way. The mass of NaCl present is calculated:

m NaCl = (10 - 6.68) g Cl2 * (2 * 58.44 g NaCl / 71 g Cl2) = 5.47 g

The percentage is calculated:

% NaCl = 5.47 * 100/10 = 54.7%

The percentage of SiO2 is calculated:

% SiO2 = 3 g * 100/10 = 30%

The percentage of NH4Cl is calculated by difference:

% NH4Cl = 100 - 54.7 - 30 = 15.3%

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.